# No Free Prune: Information-Theoretic Barriers to Pruning at Initialization

### v1

@@ Abstract
The existence of “lottery tickets” Frankle & Carbin (2018) at or near initialization raises the tantalizing question of whether large models are necessary in deep learning, or whether sparse networks can be quickly identified and trained without ever training the dense models that contain them. However, efforts to find these sparse subnetworks without training the dense model (“pruning at initialization”) have been broadly unsuccessful Frankle et al. (2020b). We put forward a theoretical explanation for this, based on the model’s effective parameter count, $p_{\text{eff}}$, given by the sum of the number of non-zero weights in the final network and the mutual information between the sparsity mask and the data. We show the Law of Robustness of Bubeck & Sellke (2023) extends to sparse networks with the usual parameter count replaced by $p_{\text{eff}}$, meaning a sparse neural network which robustly interpolates noisy data requires a heavily data-dependent mask. We posit that pruning during and after training outputs masks with higher mutual information than those produced by pruning at initialization. Thus two networks may have the same sparsities, but differ in effective parameter count based on how they were trained. This suggests that pruning near initialization may be infeasible and explains why lottery tickets exist, but cannot be found fast (i.e. without training the full network). Experiments on neural networks confirm that information gained during training may indeed affect model capacity.
@@ 1 Introduction
    At initialization, before training weights on the data Lee et al. (2018); Tanaka et al. (2020); Wang et al. (2020), often using the network’s connection structure or the loss landscape around initialization.
@@
    During training, usually in an gradual manner, starting with the dense network, training on the data, and pruning some fraction of the smallest magnitude weights, and repeating. Methods differ on what they do to weights after each prune step. The contribution of Frankle & Carbin (2018) was showing rewinding to initial weights is important for good performance at this stage of pruning. Other methods include Zhu & Gupta (2018), which iteratively prunes the smallest magnitude weights according to a predefined sparsity schedule.
@@
    After training, before inference, usually using simple but effective heuristics involving dropping the lowest magntitude weights Han et al. (2015)
Frankle & Carbin (2018) show empirically that sparse subnetworks that can train to accuracy matching or exceeding that of the full, dense model, do indeed exist at or near initialization (“matching” subnetworks). This work and the follow-ups Frankle et al. (2019, 2020a) present an algorithm to find these subnetworks called iterative magnitude pruning (IMP) with weight rewinding, dubbing such subnetworks present at initialization “lottery tickets.” In principle, this raises the enticing prospect of quickly finding these networks at initialization and training only at high sparsity, but IMP requires repeatedly training the whole model on the dataset to find these lottery tickets, defeating the original point of finding highly sparse yet trainable subnetworks.
Since then, research on “pruning at initialization” has sought to find these lottery tickets fast (i.e., without training the full model). Methods in stage (3) of the machine learning pipeline serve as benchmarks for sparsity, where those proposing pruning methods in stages (1) and (2) attempt to produce models as small as those that can be found in stage (3), where one can prune to reasonably high sparsities without compromising accuracy Frankle et al. (2020b). There remains to this day a large amount of effort to develop algorithms to prune at stage (1) Tanaka et al. (2020); Lee et al. (2018); Wang et al. (2020); Pham et al. (2022); Wang et al. (2020), but despite the diversity of techniques used, these algorithms are typically unsuccessful in finding “lottery tickets” in general settings without training. Note that our focus throughout this paper is restricted to algorithms producing matching subnetworks, as it is of course possible to prune a network at initialization if one sacrifices accuracy.
An important observation that if such an algorithm to find a matching subnetwork for a task at initialization did exist, it would suggest that the modern deep learning paradigm of training large models is misguided, as one could prune at initialization and then train the sparse subnetwork cheaply to achieve the same performance on a task. The existence of such an algorithm is also in tension with a large theoretical literature on the benefits and necessity of overparameterization Allen-Zhu et al. (2019); Du et al. (2018); Simon et al. (2023); Neyshabur (2017); Geiger et al. (2019); Bartlett et al. (2020); Montanari et al. (2024). Any theory that seeks to formalize the intractability of pruning at initialization, however, must also explain why lottery tickets can exist, but not be found efficiently (i.e., without training the full network on the data).
@@ 2 Related Work
Lottery tickets and sparsity. Pruning neural networks has a long history, from classic techniques that prune weights based on connectivity metrics involving the Jacobian and Hessian LeCun et al. (1990); Hassibi & Stork (1992) to simple but effective modern methods based on weight magnitude Han et al. (2015); Wen et al. (2016); Molchanov et al. (2016) and more recently the Lottery Ticket Hypothesis Frankle & Carbin (2018) and associated follow-up works Zhou et al. (2019); Chen et al. (2020); Frankle et al. (2020a); Gale et al. (2019); Liu et al. (2018), see Blalock et al. (2020) for a general survey. Paul et al. (2022) give a thorough loss landscape perspective on the lottery ticket hypothesis, but their work is empirical and not concerned with pruning at initialization, rather intending to illuminate the mechanism of IMP geometrically.
One important limitation of lottery tickets is that they are are subnetworks with unstructured sparsity (pruning individual weights), which is difficult to accelerate on modern hardware, whereas structured sparsity (pruning entire neurons or convolutional channels) Han et al. (2016); He et al. (2018); Zhang et al. (2020); Mao et al. (2017) is more exploitable by hardware and so often leads to more drastic performance gains despite lower levels of end-time sparsity. While we present our results to address the difficulty of finding lottery tickets, our theorems are based on notions of parameter counts which are easily adaptable to structured pruning.
Pruning at initialization. The most extensive empirical evaluation of pruning methods at initialization is Frankle et al. (2020b), which finds that the most popular methods for the task, including SNIP Lee et al. (2018), SynFlow Tanaka et al. (2020), and GraSP Wang et al. (2020), barely outperform random pruning, and are significantly beaten by even naive methods to prune after training Han et al. (2015). Their most relevant findings are:
@@
  • Frankle et al. (2020b) explicitly comment on how striking it is that methods that use such different signals (magnitudes; gradients; Hessian; data or lack thereof) end up reaching similar accuracy, behave similarly under ablations, and improve similarly when pruning after initialization. We argue there may be fundamental information-theoretic barriers causing these diverse methods to fail in very similar ways at high sparsities.
Evci et al. (2019) is another empirical work exploring the difficulty of training sparse networks, or, equivalently, pruning at initialization. Other works on efficiently finding lottery tickets include Alizadeh et al. (2022); de Jorge et al. (2021), with some new works such as Ramanujan et al. (2020); Sreenivasan et al. (2022) attempting to train a good mask at initialization by maximizing network accuracy.
Overparameterization and effective parameter count. One of the biggest surprises in modern deep learning models is that overparameterized models generalize well and don’t overfit Zhang et al. (2021). A great deal of work has gone into quantitative analysis of this mystery. We build in particular on Bubeck & Sellke (2023), where it is shown that overparameterization by a factor of the data dimension is necessary for smooth interpolation. Other releveant work includes Allen-Zhu et al. (2019), which proves that SGD can find global minima in polynomial time of number of layers and parameters if sufficiently overparameterized, and Du et al. (2018) which proves gradient descent converges in linear time to a global optimum for a two-layer ReLU network if sufficiently overparameterized. Simon et al. (2023) prove that overparameterization is necessary for near-optimal performance in random feature regression.
Mutual information and generalization bounds. Since Russo & Zou (2019) studied “bad information usage” in the setting of adaptive data analysis, much work has been done on bounding the generalization gap using the informational quantity $I(W;\mathcal{D})$ where $W$ represents the chosen hypothesis by the learning algorithm, and $D$ the dataset sampled from a data distribution Xu & Raginsky (2017); Bu et al. (2020); Asadi et al. (2018). The main takeaway from these works is that learning algorithms whose mutual information with the data is low must generalize well. From a technical viewpoint our work is closely related to these, although the motivation is different.
@@ 3 Contributions
  • We state and prove a modified version of the Law of Robustness in Bubeck & Sellke (2023) where parameter count is replaced with a data-dependent “effective parameter count” that includes both the number of parameters and the mutual information of the sparsity mask with the dataset, showing a new way in which information and parameters can be traded off. In fact, this general principle of trading off parameter count for mutual information with the data extends beyond the Law of Robustness, as we show in Appendix D.
@@ 4 Informally Stated Results & Implications
Theorem 1 (Theorem 1, Bubeck & Sellke (2023), informal).
@@
Theorem 2 (Informal, Modified Law of Robustness).
Assume the same conditions as in Theorem 1 and that $\mathcal{F}$ has the additional structure of masks, so that each hypothesis $f\in\mathcal{F}$ has parameters $(\mathbf{m},\mathbf{w})$ satisfying $\mathbf{m}_{i}=0\implies\mathbf{w}_{i}=0$ for all $i\in[p]$. Then, with high probability over sampling of the data, one has for any learning algorithm $W$ taking in data $\mathcal{D}$ and outputting function $f^{W}\in\mathcal{F}$:
[…]
@@
Overparameterization and Mutual Information. Following Bubeck & Sellke (2023), we show that for the learned function to fit below the noise level, it must correlate with the noise in the data. The key difference in our proof is a result from information theory (Lemma 1 below) which bounds this correlation by the mutual information between the learned function and the data. This bound has been used in the literature to show generalization error is small with high probability when the learned function has low mutual information with the data. Instead, we show that since the function fits below the noise level, this correlation must be large, and thus the mutual information must be large with high probability. Although classical generalization bounds recommend a low parameter count, from the perspective of the Law of Robustness (and much of modern deep learning practice), overparameterization is necessary to find a hypothesis interpolating the data smoothly. As a result, the chosen hypothesis must have high mutual information with the data for this to be possible (though it is not sufficient).
On Fitting Below Noise Level. Our results are restricted to the regime where the fitted function $f$ fits below the noise level $\sigma^{2}$ of the data, but this is a weak assumption usually satisfied in practice. For example, in computer vision, the setting in which pruning was first investigated, it is common to see networks with near-perfect test accuracy (so they have learned all the relevant signal) have end-time train loss lower than end-time test loss (so they must have memorized some noise). A typical example where such train-test loss behavior can be seen in a practical setting is in the ResNet paper He et al. (2016), and other examples include Huang et al. (2017); Lin et al. (2013). Further, Feldman (2020) suggests that this phenomenon of fitting below the noise level may be necessary for achieving optimal generalization error on real-world datasets. This is a consequence of such data often being a mixture of subpopulations where the distribution of subpopulation frequencies often follows power laws and is thus long-tailed. One can also interpret $\sigma^{2}$ as the portion of a given task that is “hard to learn” for the choice of model - that is, for a given model class, it can be seen as the residual variance conditional on a “good” feature representation, which will be larger due to portions of the target function falling outside this representation. This behavior is by now well established in several theoretical settings, see e.g. Jacot et al. (2020); Canatar et al. (2021); Ghorbani et al. (2021); Misiakiewicz & Montanari (2023). We defer to Bubeck & Sellke (2023) for further discussion.
@@ 4.1 Interpretation of $p_{\text{eff}}$
We now discuss the intuition behind the parameter $p_{\text{eff}}$ in Theorem 2. In practice, one generally attempts to prune a model to a specified sparsity level $\gamma$, so that $(1-\gamma)p$ parameters remain, with $1-\gamma$ taken to be small, often around 5% or less. Thus $\mathbb{E}[\|\mathbf{m}^{W}\|]=(1-\gamma)p$. We then compare $p_{\text{eff}}$ for the following two pruning schemes:
  • Pruning at initialization in a data agnostic manner (e.g. SynFlow Tanaka et al. (2020)). The effective parameterization is $\tilde{\Theta}((1-\gamma)p)$, reflecting that one is simply training a model of much smaller size, and hence does not obtain any benefits from the overparameterization of the initial dense model [^2] Here $\tilde{\Theta}$ hides logarithmic terms in weight size and dependence on $\delta$..
@@
    The last bound follows from Csiszár & Shields (2004) Lemma 2.2, and the total effective parameter count is this quantity combined with $\tilde{\Theta}((1-\gamma)p)$. We note that the extreme case of (1) corresponds to Eq. (2.13) in Bubeck & Sellke (2023), i.e. a worst-case bound on the effective parameter count.
@@ 5.1 Preliminaries
Isoperimetry Our results, which follow those of Bubeck & Sellke (2023), assume that the distribution of the data covariates $x_{i}$ are isoperimetric in the following sense:
Definition 1.
@@
Isoperimetry asserts that Lipschitz functions concentrate sharply around their mean, and is a ubiquitous property of “truly high-dimensional” distributions such as Gaussians. Information Theoretic Concentration Bounds
@@
To understand this result, it is helpful to consider two extremes. First if $W$ is independent from $\{X_{t}\}_{t\in T}$, then standard concentration results show (3) holds with extremely high probability without the mutual information term, so the size of $T$ is irrelevant. On the other hand, in the worst case $X_{W}$ could be the largest of $T$ subgaussian variables, causing it to scale with $\sqrt{\log|T|}$. The above result allows us to interpolate between these two regimes. We note that the overall strategy of our argument is technically very similar to that of Xu & Raginsky (2017) and related work such as Russo & Zou (2019). The conceptual difference is that we measure the information in the choice of function class rather than on the level of individual functions.

### v2

@@ Abstract
The existence of “lottery tickets” (Frankle & Carbin, 2018) at or near initialization raises the tantalizing question of whether large models are necessary in deep learning, or whether sparse networks can be quickly identified and trained without ever training the dense models that contain them. However, efforts to find these sparse subnetworks without training the dense model (“pruning at initialization”) have been broadly unsuccessful (Frankle et al., 2020b). We put forward a theoretical explanation for this, based on the model’s effective parameter count, $p_{\text{eff}}$, given by the sum of the number of non-zero weights in the final network and the mutual information between the sparsity mask and the data. We show the Law of Robustness of (Bubeck & Sellke, 2023) extends to sparse networks with the usual parameter count replaced by $p_{\text{eff}}$, meaning a sparse neural network which robustly interpolates noisy data requires a heavily data-dependent mask. We posit that pruning during and after training outputs masks with higher mutual information than those produced by pruning at initialization. Thus two networks may have the same sparsities, but differ in effective parameter count based on how they were trained. This suggests that pruning near initialization may be infeasible and explains why lottery tickets exist, but cannot be found fast (i.e. without training the full network). Experiments on neural networks confirm that information gained during training may indeed affect model capacity.
Machine Learning, ICML
@@ 1 Introduction
    At initialization, before training weights on the data (Lee et al., 2018; Tanaka et al., 2020; Wang et al., 2020), often using the network’s connection structure or the loss landscape around initialization.
@@
    During training, usually in an gradual manner, starting with the dense network, training on the data, and pruning some fraction of the smallest magnitude weights, and repeating. Methods differ on what they do to weights after each prune step. The contribution of (Frankle & Carbin, 2018) was showing rewinding to initial weights is important for good performance at this stage of pruning. Other methods include (Zhu & Gupta, 2018), which iteratively prunes the smallest magnitude weights according to a predefined sparsity schedule.
@@
    After training, before inference, usually using simple but effective heuristics involving dropping the lowest magntitude weights (Han et al., 2015)
(Frankle & Carbin, 2018) show empirically that sparse subnetworks that can train to accuracy matching or exceeding that of the full, dense model, do indeed exist at or near initialization (“matching” subnetworks). This work and the follow-ups (Frankle et al., 2019, 2020a) present an algorithm to find these subnetworks called iterative magnitude pruning (IMP) with weight rewinding, dubbing such subnetworks present at initialization “lottery tickets.” In principle, this raises the enticing prospect of quickly finding these networks at initialization and training only at high sparsity, but IMP requires repeatedly training the whole model on the dataset to find these lottery tickets, defeating the original point of finding highly sparse yet trainable subnetworks.
Since then, research on “pruning at initialization” has sought to find these lottery tickets fast (i.e., without training the full model). Methods in stage (3) of the machine learning pipeline serve as benchmarks for sparsity, where those proposing pruning methods in stages (1) and (2) attempt to produce models as small as those that can be found in stage (3), where one can prune to reasonably high sparsities without compromising accuracy (Frankle et al., 2020b). There remains to this day a large amount of effort to develop algorithms to prune at stage (1) (Tanaka et al., 2020; Lee et al., 2018; Wang et al., 2020; Pham et al., 2022; Wang et al., 2020), but despite the diversity of techniques used, these algorithms are typically unsuccessful in finding “lottery tickets” in general settings without training. Note that our focus throughout this paper is restricted to algorithms producing matching subnetworks, as it is of course possible to prune a network at initialization if one sacrifices accuracy.
An important observation that if such an algorithm to find a matching subnetwork for a task at initialization did exist, it would suggest that the modern deep learning paradigm of training large models is misguided, as one could prune at initialization and then train the sparse subnetwork cheaply to achieve the same performance on a task. The existence of such an algorithm is also in tension with a large theoretical literature on the benefits and necessity of overparameterization (Allen-Zhu et al., 2019; Du et al., 2018; Simon et al., 2023; Neyshabur, 2017; Geiger et al., 2019; Bartlett et al., 2020; Montanari et al., 2024). Any theory that seeks to formalize the intractability of pruning at initialization, however, must also explain why lottery tickets can exist, but not be found efficiently (i.e., without training the full network on the data).
@@ 2 Related Work
Lottery tickets and sparsity. Pruning neural networks has a long history, from classic techniques that prune weights based on connectivity metrics involving the Jacobian and Hessian (LeCun et al., 1990; Hassibi & Stork, 1992) to simple but effective modern methods based on weight magnitude (Han et al., 2015; Wen et al., 2016; Molchanov et al., 2016) and more recently the Lottery Ticket Hypothesis (Frankle & Carbin, 2018) and associated follow-up works (Zhou et al., 2019; Chen et al., 2020; Frankle et al., 2020a; Gale et al., 2019; Liu et al., 2018), see (Blalock et al., 2020) for a general survey. (Paul et al., 2022) give a thorough loss landscape perspective on the lottery ticket hypothesis, but their work is empirical and not concerned with pruning at initialization, rather intending to illuminate the mechanism of IMP geometrically.
One important limitation of lottery tickets is that they are are subnetworks with unstructured sparsity (pruning individual weights), which is difficult to accelerate on modern hardware, whereas structured sparsity (pruning entire neurons or convolutional channels) (Han et al., 2016; He et al., 2018; Zhang et al., 2020; Mao et al., 2017) is more exploitable by hardware and so often leads to more drastic performance gains despite lower levels of end-time sparsity. While we present our results to address the difficulty of finding lottery tickets, our theorems are based on notions of parameter counts which are easily adaptable to structured pruning.
Pruning at initialization. The most extensive empirical evaluation of pruning methods at initialization is (Frankle et al., 2020b), which finds that the most popular methods for the task, including SNIP (Lee et al., 2018), SynFlow (Tanaka et al., 2020), and GraSP (Wang et al., 2020), barely outperform random pruning, and are significantly beaten by even naive methods to prune after training (Han et al., 2015). Their most relevant findings are:
@@
  • (Frankle et al., 2020b) explicitly comment on how striking it is that methods that use such different signals (magnitudes; gradients; Hessian; data or lack thereof) end up reaching similar accuracy, behave similarly under ablations, and improve similarly when pruning after initialization. We argue there may be fundamental information-theoretic barriers causing these diverse methods to fail in very similar ways at high sparsities.
(Evci et al., 2019) is another empirical work exploring the difficulty of training sparse networks, or, equivalently, pruning at initialization. Other works on efficiently finding lottery tickets include (Alizadeh et al., 2022; de Jorge et al., 2021), with some new works such as (Ramanujan et al., 2020; Sreenivasan et al., 2022) attempting to train a good mask at initialization by maximizing network accuracy.
Overparameterization and effective parameter count. One of the biggest surprises in modern deep learning models is that overparameterized models generalize well and don’t overfit (Zhang et al., 2021). A great deal of work has gone into quantitative analysis of this mystery. We build in particular on (Bubeck & Sellke, 2023), where it is shown that overparameterization by a factor of the data dimension is necessary for smooth interpolation. Other releveant work includes (Allen-Zhu et al., 2019), which proves that SGD can find global minima in polynomial time of number of layers and parameters if sufficiently overparameterized, and (Du et al., 2018) which proves gradient descent converges in linear time to a global optimum for a two-layer ReLU network if sufficiently overparameterized. (Simon et al., 2023) prove that overparameterization is necessary for near-optimal performance in random feature regression.
Mutual information and generalization bounds. Since (Russo & Zou, 2019) studied “bad information usage” in the setting of adaptive data analysis, much work has been done on bounding the generalization gap using the informational quantity $I(W;\mathcal{D})$ where $W$ represents the chosen hypothesis by the learning algorithm, and $D$ the dataset sampled from a data distribution (Xu & Raginsky, 2017; Bu et al., 2020; Asadi et al., 2018). The main takeaway from these works is that learning algorithms whose mutual information with the data is low must generalize well. From a technical viewpoint our work is closely related to these, although the motivation is different.
@@ 3 Contributions
  • We state and prove a modified version of the Law of Robustness in (Bubeck & Sellke, 2023) where parameter count is replaced with a data-dependent “effective parameter count” that includes both the number of parameters and the mutual information of the sparsity mask with the dataset, showing a new way in which information and parameters can be traded off. In fact, this general principle of trading off parameter count for mutual information with the data extends beyond the Law of Robustness, as we show in Appendix C.
@@ 4 Informally Stated Results & Implications
Theorem 4.1 (Theorem 1, (Bubeck & Sellke, 2023), informal).
@@
Theorem 4.2 (Informal, Modified Law of Robustness).
Assume the same conditions as in Theorem 4.1 and that $\mathcal{F}$ has the additional structure of masks, so that each hypothesis $f\in\mathcal{F}$ has parameters $(\mathbf{m},\mathbf{w})$ satisfying $\mathbf{m}_{i}=0\implies\mathbf{w}_{i}=0$ for all $i\in[p]$. Then, with high probability over sampling of the data, one has for any learning algorithm $W$ taking in data $\mathcal{D}$ and outputting function $f^{W}\in\mathcal{F}$:
$\frac{1}{n}\sum_{i=1}^{n}(f^{W}(x_{i})-y_{i})^{2}\leq\sigma^{2}-\epsilon\\ \implies\mathrm{Lip}(f)\geq\widetilde{\Omega}\left(\epsilon\sqrt{\frac{nd}{p_{\text{eff}}}}\right),$ (1)
where $p_{\text{eff}}=\tilde{\Theta}\left(I(\mathbf{m}^{W};\mathcal{D})+\mathbb{E}[\|\mathbf{m}\|_{1}]\right)$.
@@
We give an example which saturates the bound presented above in Appendix C.1.1.
Overparameterization and Mutual Information. Following (Bubeck & Sellke, 2023), we show that for the learned function to fit below the noise level, it must correlate with the noise in the data. The key difference in our proof is a result from information theory (Lemma 5.2 below) which bounds this correlation by the mutual information between the learned function and the data. This bound has been used in the literature to show generalization error is small with high probability when the learned function has low mutual information with the data. Instead, we show that since the function fits below the noise level, this correlation must be large, and thus the mutual information must be large with high probability. Although classical generalization bounds recommend a low parameter count, from the perspective of the Law of Robustness (and much of modern deep learning practice), overparameterization is necessary to find a hypothesis interpolating the data smoothly. As a result, the chosen hypothesis must have high mutual information with the data for this to be possible (though it is not sufficient).
On Fitting Below Noise Level. Our results are restricted to the regime where the fitted function $f$ fits below the noise level $\sigma^{2}$ of the data, but this is a weak assumption usually satisfied in practice. For example, in computer vision, the setting in which pruning was first investigated, it is common to see networks with near-perfect test accuracy (so they have learned all the relevant signal) have end-time train loss lower than end-time test loss (so they must have memorized some noise). A typical example where such train-test loss behavior can be seen in a practical setting is in the ResNet paper (He et al., 2016), and other examples include (Huang et al., 2017; Lin et al., 2013). Further, (Feldman, 2020) suggests that this phenomenon of fitting below the noise level may be necessary for achieving optimal generalization error on real-world datasets. This is a consequence of such data often being a mixture of subpopulations where the distribution of subpopulation frequencies often follows power laws and is thus long-tailed. One can also interpret $\sigma^{2}$ as the portion of a given task that is “hard to learn” for the choice of model - that is, for a given model class, it can be seen as the residual variance conditional on a “good” feature representation, which will be larger due to portions of the target function falling outside this representation. This behavior is by now well established in several theoretical settings, see e.g. (Jacot et al., 2020; Canatar et al., 2021; Ghorbani et al., 2021; Misiakiewicz & Montanari, 2023). We defer to (Bubeck & Sellke, 2023) for further discussion.
@@ 4.1 Interpretation of $p_{\text{eff}}$
We now discuss the intuition behind the parameter $p_{\text{eff}}$ in Theorem 4.2. In practice, one generally attempts to prune a model to a specified sparsity level $\gamma$, so that $(1-\gamma)p$ parameters remain, with $1-\gamma$ taken to be small, often around 5% or less. Thus $\mathbb{E}[\|\mathbf{m}^{W}\|]=(1-\gamma)p$. We then compare $p_{\text{eff}}$ for the following two pruning schemes:
  • Pruning at initialization in a data agnostic manner (e.g. SynFlow (Tanaka et al., 2020)). The effective parameterization is $\tilde{\Theta}((1-\gamma)p)$, reflecting that one is simply training a model of much smaller size, and hence does not obtain any benefits from the overparameterization of the initial dense model [^2] Here $\tilde{\Theta}$ hides logarithmic terms in weight size and dependence on $\delta$..
@@
    The last bound follows from (Csiszár & Shields, 2004) Lemma 2.2, and the total effective parameter count is this quantity combined with $\tilde{\Theta}((1-\gamma)p)$. We note that the extreme case of (2) corresponds to Eq. (2.13) in (Bubeck & Sellke, 2023), i.e. a worst-case bound on the effective parameter count.
@@ 5.1 Preliminaries
Isoperimetry Our results, which follow those of (Bubeck & Sellke, 2023), assume that the distribution of the data covariates $x_{i}$ are isoperimetric in the following sense:
Definition 5.1.
@@
Isoperimetry asserts that Lipschitz functions concentrate sharply around their mean, and is a ubiquitous property of “truly high-dimensional” distributions such as Gaussians.
Information Theoretic Concentration Bounds
@@
To understand this result, it is helpful to consider two extremes. First if $W$ is independent from $\{X_{t}\}_{t\in T}$, then standard concentration results show (4) holds with extremely high probability without the mutual information term, so the size of $T$ is irrelevant. On the other hand, in the worst case $X_{W}$ could be the largest of $T$ subgaussian variables, causing it to scale with $\sqrt{\log|T|}$. The above result allows us to interpolate between these two regimes. We note that the overall strategy of our argument is technically very similar to that of (Xu & Raginsky, 2017) and related work such as (Russo & Zou, 2019). The conceptual difference is that we measure the information in the choice of function class rather than on the level of individual functions.